A population consists of the five measurements 2, 6, 8, 0, and 1 with a sample size 4. Find the variance of the sample means


1
Expert's answer
2021-05-21T02:18:13-0400

Mean



"\mu=\dfrac{2+6+8+0+1}{5}=3.4"

Variance


"\sigma^2=\dfrac{1}{5}\big((2-3.4)^2+(6-2.4)^2+(8-3.4)^2""(0-3.4)^2+(1-3.4)^2\big)=9.44"


Standard deviation



"\sigma=\sqrt{\sigma^2}=\sqrt{9.44}\approx3.0725"




 We have population values "2,6,8,0,1" population size "N=5" and sample size "n=4." Thus, the number of possible samples which can be drawn without replacement is


"\dbinom{N}{n}=\dbinom{5}{4}=5""\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & Sample & Sample \ mean \\
 No. & values & (\bar{X}) \\ \hline
 1 & 2,6, 8, 0 & 4 \\
 \hdashline
 2 & 2,6,8,1 & 4.25 \\
 \hdashline
 3 & 2,6,0,1 & 2.25 \\
 \hdashline
 4 & 2,8,0,1 & 2.75 \\
 \hdashline
 5 & 6,8,0,1 & 3.75 \\
 \hline
\end{array}"





"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f & f(\bar{X}) & \bar{X}f(\bar{X})& \bar{X}^2f(\bar{X}) \\ \hline
 9/4 & 1& 1/5 & 9/20 & 81/80 \\
 \hdashline
 11/4 & 1 & 1/5 & 11/20 & 121/80 \\
 \hdashline
 15/4 & 1 & 1/5 & 15/20 & 225/80 \\
 \hdashline
 4 & 1 & 1/5 & 16/20 & 256/80 \\
 \hdashline
 17/4 & 1& 1/5 & 17/20 & 289/80 \\
 \hdashline
 Total & 5 & 1 & 68/20 & 972/80 \\ \hline
\end{array}"




"E(\bar{X})=\sum\bar{X}f(\bar{X})=\dfrac{68}{20}=3.4"

The mean of the sampling distribution of the sample means is equal to the

the mean of the population.



"E(\bar{X})=3.4=\mu"




"Var(\bar{X})=\sum\bar{X}^2f(\bar{X})-(\sum\bar{X}f(\bar{X}))^2"




"=\dfrac{972}{80}-(\dfrac{68}{20})^2=\dfrac{472}{800}=\dfrac{59}{100}=0.59"




"\sqrt{Var(\bar{X})}=\sqrt{\dfrac{59}{100}}\approx0.7681"

Verification:


"Var(\bar{X})=\dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})=\dfrac{9.44}{4}(\dfrac{5-4}{5-1})""=0.59, True"